OpenStudy (anonymous):

determine the following intergral:(ax + b)^(3/2) dx

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

ax^(a-1)?

OpenStudy (anonymous):

no i am not familiar with it

OpenStudy (anonymous):

sorry, i thought , it differentiation....

sam (.sam.):

try let u=ax+b du=adx \[\frac{1}{a}\int\limits u^{3/2} \, du\]

sam (.sam.):

then just integrate and sub back in

OpenStudy (anonymous):

int of x^n dx=x^(n+1)/n+1

OpenStudy (anonymous):

how does one intergrate though

OpenStudy (anonymous):

Follow @Sam suggestion. He has given u some clue.

OpenStudy (anonymous):

Try first

sam (.sam.):

use this

OpenStudy (kropot72):

\[\int\limits_{}^{}(ax+b)^{\frac{3}{2}}dx=\frac{2(ax+b)^{\frac{5}{2}}}{5a}\]

OpenStudy (kropot72):

Plus C of course.

OpenStudy (kropot72):

This was done by inspection using the standard rule for integration ahd then differentiating to check and adjust the result.